# Joint Denoising and Fusion with Short- and Long-exposure Raw Pairs

### v4

@@ 3.2.2 Data Processing
Fig. 2 shows the proposed data formation pipeline. Here’s an in-depth breakdown of the data formation pipeline:
Selection of exposure time ratio and initial adjustment. Commencing the pipeline, we pack two clean Bayer raw images. We randomly select an exposure time ratio $r$ from {4, 8, 16} at this stage.  [^1] It’s important to note that exposure ratios may not necessarily be multiples of 2. The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, multiply by the maximum pixel value (Max: $2^{12}\times r$). This operation yields $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{2}$ correspondingly.
Long-exposure image simulation ($I_{l}$). Moving forward, we divide $I^{{}^{\prime}}_{1}$ by 1 and then clip the pixel values to a range of 0 to 4095 (12 bit). This operation is equivalent to adjusting the Digital Gain ($K{\textrm{d}}$), achieving an outcome comparable to altering the exposure time. Subsequently, we synthesize the noisy image by long exposure image noise model measured by Sony IMX327 HDR sensor and normalize it. This process simulates a long-exposure noisy image ($I_{l}$) reaching saturation signal level (full well capacity) and the inherent noise generated by the image sensor, which preserves dark detail while losing details in brighter areas.
Short-exposure image simulation ($I_{l}$). Simultaneously, we divide $I^{{}^{\prime}}_{2}$ by $r$ and clip the pixel values to a range of 0 to 4095. (Equivalent to adjusting the Digital Gain). Similar to the previous step, synthesize the short-exposure noisy image by the short-exposure image noise model measured by the Sony IMX327 HDR sensor, and normalization follows suit. This procedure simulates a practical short-exposure image ($I_{s}$). This image retains detail in highlighted portions but loses darker information due to noise interference.
Ground truth image ($I_{gt}$). Additionally, we normalize $I^{{}^{\prime}}_{2}$ to obtain the ground truth image $I_{gt}$. The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions compared to $I_{s}$. As a result, our data formation pipeline efficiently generates an extensive array of LDR-HDR data pairs.
@@
The data formation pipeline facilitates the generation of rich LDR-HDR pairs, which is integral to our research and exploration.
Figure 2: The raw LDR-HDR pair formation pipeline. Two clean high-bit raw images, $I_{1}$ and $I_{2}$, have been processed through black-level correction, normalization, and BM3D denoising.
@@ 3.2.3 RealRaw-HDR Dataset
Our dataset is meticulously crafted for dual-exposure HDR sensors, supporting mainstream sensors, including Sony IMX327, IMX385, IMX585, and OV OS05B. To the best of our knowledge, there is an absence of a raw HDR dataset explicitly tailored for these HDR sensors. Furthermore, it is challenging to construct large-scale, high-quality raw HDR datasets from multi-exposed frames. Traditional methods of collecting HDR datasets often lead to misalignment issues, yielding fused HDR images of subpar quality; previous data collection methods necessitate subjects to remain static while capturing multiple exposures, a time-consuming and labor-intensive process. In contrast, our proposed raw LDR-HDR pair formation pipeline is efficient and user-friendly, enabling the effortless creation of a multitude of high-quality data pairs.
Through this pipeline, we gather $240$ pairs of $8192\times 6192$ high-resolution raw LDR-HDR image pairs. Fig. 3 shows an example of two generated LDR-HDR pairs with different exposure ratios. Additionally, by attaching an ISP pipeline to the end of our pipeline, we can create an sRGB-based HDR training dataset. In Tab. 1, we compare the statistics of our dataset with those of other existing HDR datasets.  [^2] If not specified, all raw images in this paper have been processed by a fixed ISP to obtain sRGB versions for visualization purposes.
Table 1: Comparison with existing HDR datasets. ×$\times 3$ indicates that the same set of images can be expanded with three exposure ratios. Theoretically, Our dataset can be expanded $N$ times by the raw LDR-HDR pair formation pipeline, N denotes the number of covered exposure ratios.
@@
Kalanatri [16] 74 $1490\times 989$ sRGB {4, 8, 16}
Chen [5] 144 $4096\times 2168$ raw, sRGB {4, 8, 16}
Ours $240\times 3$ $8192\times 6192$ raw, sRGB {4, 8, 16}
4 The Proposed RepUNet Network
HDR reconstruction plays a vital role in various applications, such as mobile photography, high-definition displays, and virtual reality, where lightweight and efficient algorithms are highly demanded due to resource limitations. Previous learning-based HDR methods [22, 21, 16] often rely on large and complex models, making them impractical for real-world scenarios. Therefore, we propose a UNet-like lightweight base model, DualUNet, to address this challenge and employ reparameterization techniques to improve its representation capability while balancing storage and computation cost and reconstruction accuracy. We first present the DualUNet and then develop our RepUNet by reparameterization, followed by the design of the loss function.
4.1 Base Model of DualUNet
To ensure high inference speed and cross-device deployment on commodity mobile devices, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a UNet-like network as the base model, namely DualUNet. The overall architecture of the DualUNet is shown in Fig. 4(a). We do not use the computationally demanding explicit alignment in our HDR network for lightweight design. We introduce an alignment-free and motion-aware short-exposure-first selection loss (in Sec. 4.3) that enables training with unaligned pairs to compensate for the absence of alignment modules.
We first use the pixel unshuffle operation [10] to transfer the input raw images $I^{\textquoteright}_{1}$ and $I^{\textquoteright}_{s}$ from $C\times H\times W$ to $4C\times\frac{1}{2}H\times\frac{1}{2}W$ to extract large-scale contextual information while maintaining low computational complexity. Notably, short-exposure images have less motion blur distortion but severe noise in dark areas. Conversely, long-exposure images generally contain more information with a higher signal-to-noise ratio but may suffer from significant motion blur and overexposure.
To better leverage the complementary information from the two types of images, we introduce two distinct sub-encoders: Encoder-S and Encoder-L. Encoder-S extracts features from the short-exposure image, serving as foundational reference features. In parallel, Encoder-L extracts features from the long-exposure image, offering supplementary features. We fuse the multi-scale features using addition or concatenation operations at each feature level to facilitate complementary features to learn motion information from the reference features.
Each Encoder-S consists of a pixel unshuffle $\downarrow 2$ downsampling operation, a $3\times 3$ convolution layer, and a ReLU activation ([Down-Conv-ReLU]). At the same time, the Encoder-L has the structure of [Down-Conv-ReLU-Conv-ReLU]. The pixel unshuffle operation can avoid information loss introduced by conventional downsampling/pooling operations. Finally, the fused features are fed into the decoder to generate the final HDR output.
Refer to caption (a) DualUNet
Refer to caption (b) Topology Convolution Block (TCB)
@@ 4.1 Base Model of DualUNet
Figure 4: Illustration of (a) DualUNet and (b) Topology Convolution Block (TCB). In the training phase, the TCB module replaces the normal convolution, which is merged back into a normal 3 × 3 convolution layer in the inference phase.
4.2 RepUNet Network
The representation capability of the plain DualUNet for HDR imaging is significantly limited, as shown in Tab. 6, leading to unsatisfactory performance. To address this concern, we introduce a flexible re-parameterizable module called the Topological Convolution Block (TCB) and replace the normal convolution in DualUNet with the TCB module during the training phase. Notably, the last layer of DualUNet is still a normal convolution.
Although the reparameterization has achieved promising results on other tasks [9, 4, 7, 46], its application in reparameterization blocks for low-level visual tasks remains scarce. We integrate previous reparameterization blocks and flexibly design reparameterization blocks with six operator combinations to enhance the feature representation of the model for HDR and denoising tasks. The TCB consists of several fundamental units: (1) a standard $3\times 3$ convolution for a solid foundation; (2) extending and squeezing convolution to enhance feature expressiveness, (3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges; (4) a jump connection to avoid gradient vanishing or exploding, and (5) a $1\times 1$ convolution to encourage information exchange between channels. It’s paramount to underscore that we exclusively employ the TCB with the Laplacian operator within the decoder. This selective approach is grounded in Laplacian operator effectiveness for noise-free images, underpinning its application to enhance feature representation in contexts devoid of noise.
To achieve an efficient HDR network that meets the stipulated design prerequisites of low computational complexity and streamlined hardware device deployment, we simplify the TCB reparameterization into a single $3\times 3$ convolution after training. Following previous works [8, 7, 46], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolution into a normal $3\times 3$ convolution. Additionally, we combine the Sobel and Laplacian operators into a special $3\times 3$ convolution with a fixed convolution kernel. The $1\times 1$ convolution is achieved by padding the convolution kernel with zeros. As a result, TCB can be transformed into a single $3\times 3$ convolution for efficient implementation during the inference stage. By utilizing TCB, we achieve superior HDR results with improved efficiency.
4.3 Loss Functions
Alignment-free and motion-aware short-exposure-first selection loss. In fused-based HDR imaging, eliminating ghosting caused by motion inconsistencies between short and long frames is one of the most challenging issues. Previous work [16, 37] commonly employs optical flow, attention mechanisms, and other methods to establish pixel correspondences between long and short frames. The objective is to suppress ghosting by designing more elaborate fusion strategies. Yet, these motion estimation and alignment methods are often the most computationally intensive components and cannot be accommodated by the current level of hardware design. To ensure fast inference and easy hardware deployment, we abandon explicit motion alignment schemes and instead explore a more efficient alternative. Consequently, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate the ghost artifacts.
Unlike other tasks involving image alignment, such as video motion estimation and stereo matching, long and short-exposure image fusion in HDR reconstruction does not necessarily require pixel-level correspondence in real-world scenes. The reason is that it is challenging to recover sharp object edges due to motion blur. In contrast, short-exposure images exhibit less motion blur distortion. Therefore, ghost artifacts can be suppressed by simply detecting motion regions in long-frame images through some mechanism, discarding these pixels during the fusion process, and relying solely on the corresponding regions in short-exposure images as the exclusive information source for fusion. Based on the same consideration, overexposed regions in the long-exposure image should likewise be discarded from the fusion process. Therefore, in practical engineering applications, short exposures are often used as reference images.
Based on the above analysis, we devise the following strategies for double-exposure HDR fusion: in scenarios involving motion or overexposure within the fused region, we prefer to select the short-exposure image that contains more information; when the signal-to-noise ratio of the short frame is too low, we prefer to select the long-exposure image that contains more information. It is worth noting that our strategy is inclined to address motion artifacts with a higher priority than low signal-to-noise ratios. Consequently, when both ghost artifacts and low signal-to-noise ratios are concurrently evident, the short-exposure image is chosen to prioritize the motion artifacts.
Figure 5: An illustrative sample of data construction for the proposed pseudo location-based motion alignment loss.
@@ 4.3 Loss Functions
Note that after the luminance alignment, the long and short frames already have similar luminance levels, and the regions with remarkable luminance differences are mainly caused by motion and overexposure. Therefore, we expect the network to perceive the motion and overexposed regions by focusing on the luminance gap between the long and short exposures, thereby realizing the short-frame prioritization strategy mentioned above. For this purpose, we design an alignment-free and motion-aware short-exposure-first selection loss function.
$L_{\textrm{AMSS}}=1-MSSSIM(\widetilde{I}^{out}\odot M,{I}^{gt}\odot M),$ (6)
where $\odot$ denotes the point-wise multiplication, and $M$ is a binary mask with “1” for motion regions in long frames, and “0” otherwise. Given a mask $M$ indicating motion and overexposed regions, the above loss formula implements a strategy that encourages short-frame prioritization.
To avoid explicit motion alignment, we adopt the synthetic motion to construct the mask $M$ in the training quadruplets {$I^{{}^{\prime}}_{l}$, $I^{{}^{\prime}}_{s}$, $M$, $I^{{}^{\prime}}_{gt}$. Specifically, as shown in Fig. 5, for each patch of an image, we randomly select a rectangle (40 to 60-pixel range) of random length and width from the long-exposure image and then move and overlap it to a random location in the range of -30 to 30 relative to the patch. The patch regions before and after the movement are labeled as 1s in the mask $M$. By introducing the mask $M$ to guide the network to focus on moving and overexposed regions, the model effectively prioritizes the short-exposure information over the long-exposed counterpart within these regions. The masks are used only in training, and the inference stage inputs only short and long exposure frames.
Reconstruction loss. For saturated areas, the L2 loss punishes any deviation of pixel values from the ground truth. This allows the model to select short-exposure information in over-exposed areas. To achieve the best HDR reconstruction results, we employ the MS-SSIM loss guide model, which learns short-exposure image information for global motion. By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
$L_{\textrm{pix}}=\|\widetilde{I}^{out}-{I}^{gt}\|_{2},$ (7)
$L_{\textrm{ssim}}=1-MSSSIM(\widetilde{I}^{out},{I}^{gt}).$ (8)
@@
Table 2: Performance comparison of different HDR models on our synthetic dataset. Params and FLOPs represent the total number of network parameters and floating-point operations. The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images. Metrics with $\uparrow$ and $\downarrow$ denote higher better and lower better, respectively. The best and second-best performances are in bold and underlined, respectively. "-" indicates the result is not available.
Methods FLOPs Params Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16

### v5

@@ 3.2.2 Data Processing
Based on the digital camera imaging theory, we utilize two raw images ($I_{1}$ and $I_{2}$ have relative motion and are noise-free) to simulate short- and long-exposure images and construct the corresponding ground truths based on the principles of HDR synthesis. The Fig. 1 shows the proposed data formation pipeline.
Selection of exposure time ratio and initial adjustment. Commencing the pipeline, we pack two clean Bayer raw images. We select an exposure time ratio $r$ from {4, 8, 16} at this stage. The two normalized raw images, denoted as $I_{1}$ and $I_{2}$, multiply by the maximum pixel value (Max: $2^{12}\times r$). This operation yields $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{2}$ correspondingly.
Long-exposure image simulation ($I_{l}$). Moving forward, we divide $I^{{}^{\prime}}_{1}$ by 1 and then clip the pixel values to a range of 0 to 4095 (12 bit). This operation is equivalent to adjusting the sensor gain ($K{\textrm{d}}$), achieving an outcome comparable to altering the exposure time (as shown in Fig. 2). Then, we add noise to create the corresponding noisy long-exposure raw images. This process simulates a long-exposure noisy image ($I_{l}$) reaching saturation signal level (full well capacity) and the inherent noise generated by the image sensor, which preserves dark detail while losing details in brighter areas.
Short-exposure image simulation ($I_{l}$). Simultaneously, we divide $I^{{}^{\prime}}_{2}$ by $r$ and clip the pixel values to a range of 0 to 4095. (Equivalent to adjusting the sensor gain). Similar to the previous step, we add noise to create the corresponding noisy short-exposure raw images. This procedure simulates a practical short-exposure image ($I_{s}$). This image retains detail in highlighted portions but loses darker information due to noise interference.
Ground truth image ($I_{gt}$). HDR aims to recover detailed information from LDR images in both brighter and dark areas. Therefore, for a dual-exposure HDR sensor, we aim to recover the darkest areas of the short-exposure image from the long-exposure image. Thus, based on this principle, we normalize $I^{{}^{\prime}}_{2}$ to obtain the ground truth image $I_{gt}$. The $I_{gt}$ contains more information on bright regions than $I_{l}$; $I_{gt}$ has a higher signal-to-noise ratio in dark regions than $I_{s}$. As a result, our data formation pipeline efficiently generates an extensive array of LDR-HDR data pairs.
@@
Our degraded dataset follows the principle of HDR synthesis—namely, the principle of maximum signal-to-noise ratio. In the darkest region, we select long-exposure images; in the brightest region, we select short-exposure images.
@@ 3.2.3 RealRaw-HDR Dataset
Our dataset is meticulously crafted for dual-exposure HDR sensors, supporting mainstream sensors, including Sony IMX327, IMX385, IMX585, and OV OS05B. To the best of our knowledge, there is an absence of a raw HDR dataset explicitly tailored for these HDR sensors. Our proposed data formation pipeline is efficient and user-friendly, enabling the creation of many high-quality data pairs effortlessly. We gather $240$ pairs of $8192\times 6192$ high-resolution raw image pairs and expand to 720 pairs. Fig. 2 shows an example of two generated LDR-HDR pairs with different exposure ratios. Additionally, by attaching an ISP pipeline to the end of our pipeline, we can create an sRGB-based HDR training dataset. In Tab. 1, we compare the statistics of our dataset with those of other existing HDR datasets. In this paper, all raw images have been processed with a fixed ISP, and HDR images are processed with the same tone mapping operator to obtain the sRGB version for visualization.
Table 1: The statistics comparison between Kalanatri [14], Chen [4] and our RealRaw-HDR dataset.
@@
Kalanatri [14] 74 $1490\times 989$ sRGB 4 &\&& 8 &\&& 16
Chen [4] 144 $4096\times 2168$ raw, sRGB 4 &\&& 8 &\&& 16
Ours 720 $8192\times 6192$ raw, sRGB 4-16
3.2.4 Effectiveness of the Data Formation Pipeline
The proposed pipeline for generating HDR data is efficient and user-friendly, allowing easy generation of numerous high-quality data pairs. Although our ground truth is derived from a single image, it contains a wide range of information characteristics of HDR images. Firstly, the raw images $I_{1}$ and $I_{2}$ are captured by a high dynamic range camera and contain high dynamic range. On the other hand, there is a significant difference in the signal-to-noise ratio between HDR images and LDR images. The exposure-aligned long-exposure images differ from the short-exposure images only in the dark and overexposed regions, in addition to the noise difference. Therefore, $I_{l}$ and $I_{s}$ have all the characteristics of real-world long and short-exposure images, and $I_{gt}$ contains a wide range of informative features of real HDR images.
4 Methodology
4.1 Overview
HDR reconstruction plays a vital role in various applications, such as mobile photography, high-definition displays, and virtual reality, where lightweight and efficient algorithms are highly demanded due to resource limitations. Previous learning-based HDR methods [21, 20, 33] often rely on large and complex models, making them impractical for real-world scenarios. On the other hand, unlike GPU servers, existing optimized manipulations for mobile devices are quite limited, especially on computationally limited devices. Unsupported operators have to be processed on the CPU, which not only very low processing speed, but also introduces additional MACs. Therefore, we first design a neat UNet with mobile-friendly operations as the base model, then propose a re-parameterizable Topological Convolution Block to improve HDR performance. For lightweight design, we do not use the computationally demanding explicit alignment in our HDR network. To compensate for the absence of alignment modules, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss (in Sec. 4.5) that enables training with unaligned pairs.
4.2 Base Model
To ensure high inference speed and cross-device deployment on commodity mobile devices, we carefully consider the limited computation and memory resources on mobile devices and deliberately choose a neat UNet consisting of the most basic operations as the base model. The overall architecture of the base model is shown in Fig. 3(a), maximizing the use of the dual-exposure HDR sensor imaging characteristics. We introduce two distinct sub-encoders based on the differences in long- and short-exposure image features: Encoder-S and Encoder-L. Encoder-S extracts features from the short-exposure image, serving as reference features. In parallel, Encoder-L extracts features from the long-exposure image, offering supplementary features.
Considering the limited bandwidth, we first employ the pixel unshuffle operation [8] to transfer the input raw images $I^{{}^{\prime}}_{1}$ and $I^{{}^{\prime}}_{s}$ from $C\times H\times W$ to $4C\times\frac{H}{2}\times\frac{W}{2}$ to extract multi-scale contextual information while keeping the MAC of our model as low as possible. Subsequently, reference and complementary features are extracted by different numbers of normal convolutions, respectively. To be more specific, each layer of Encoder-S consists of a pixel unshuffle $\downarrow 2$ downsampling operation, a $3\times 3$ convolution layer, and a ReLU activation ([Down-Conv-ReLU]). At the same time, each layer of Encoder-L has the structure of [Down-Conv-ReLU-Conv-ReLU]. To promote the complementary features to learn the relative motion from the reference features, we feed the reference features of each layer to the next layer by the addition of the reference features with the complementary features. Finally, we concatenate the reference features with the complementary features and feed them to the decoder. The decoder only contains 5 normal convolutions and 3 upsampling operators. By delicate design, the proposed base model is well-suitable for mobile scenarios with high efficiency and flexibility. The network design with low MAC allows for ultra-fast inference on mobile devices, and the basic operation makes cross-device deployment easier.
(a) Base Model
(b) Topology Convolution Block (TCB)
@@ 4.2 Base Model
Figure 3: Illustration of (a) Base Model and (b) Topology Convolution Block (TCB). In the training phase, the TCB employs multiple branches, which can be merged into one normal convolution layer in the inference stage.
@@
4.3 Topological Convolution Block
Although the plain base model is efficient, its HDR performance is less satisfactory compared to those complicated models, as shown in Tab. 5. We thus employ the re-parameterization technique to enrich the representation capability of the base model. The reparameterization has achieved promising results on other tasks [7, 3, 5, 44]. We design a flexible re-parameterizable module called the Topological Convolution Block (TCB)), which can more effectively extract edge and texture information for the HDR task. As shown in Fig. 3(b), the TCB consists of several fundamental units: (1) A standard $3\times 3$ convolution for a solid foundation. The standard convolution is denoted as:
$F_{n}=W_{n}*X+B_{n},$ (6)
where $F_{n}$, $X$, $W_{n}$, and $B_{n}$ represent the output feature, input feature, weights, and bias of the standard convolution, respectively.
(2) Extending and squeezing convolution to enhance feature expressiveness, the expanding and squeezing feature is extracted as:
$F_{es}=W_{s}*(W_{e}*X+B_{e})+B_{s},$ (7)
where $W_{e}$, $B_{e}$ and $W_{s}$, $B_{s}$ are the $1\times 1$ expanding and $3\times 3$ squeezing convolutions weights, bias, respectively.
(3) Sobel and Laplacian operators for extracting first and second-order spatial derivatives to identify edges, i.e., using a predetermined convolution kernel to process the edge information. Denote by $D_{x}$ and $D_{y}$ the horizontal and vertical Sobel filters, $D_{lap}$ is the Laplacian filter.
$D_{x}=\begin{bmatrix}+1&0&-1\\ +2&0&-2\\ +1&0&-1\end{bmatrix},$ (8)
$D_{y}=\begin{bmatrix}+1&+2&+1\\ 0&0&0\\ -1&-2&-1\end{bmatrix},$ (9)
$D_{lap}=\begin{bmatrix}0&+1&0\\ +1&-4&+1\\ 0&+1&0\end{bmatrix}$ (10)
The combined edge information is extracted by:
$F_{edge}=F_{Dx}+F_{Dy}+F_{lap},$ (11)
where $F_{Dx}$, $F_{Dy}$, and $F_{lap}$ represent the horizontal, vertical, 2nd-order edge information, respectively.
(4) A $1\times 1$ convolution to encourage information exchange between channels, denoted as:
$F_{c}=W_{c}*X+B_{c},$ (12)
where $F_{c}$, $W_{c}$, and $B_{c}$ represent the output feature, weights, and bias of the $1\times 1$ convolution, respectively.
(5) A jump connection to avoid gradient vanishing or exploding, denoted as:
$F_{j}=X$ (13)
The output of the TCB in the combination of the five components:
$F_{TCB}=F_{n}+F_{es}+F_{edge}+F_{c}+F_{j}$ (14)
The combined feature map is then fed into a non-linear activation layer. PReLU is employed in our experiments. It is paramount to underscore that we exclusively employ the TCB with the Laplacian operator within the decoder. This selective approach is grounded in Laplacian operator effectiveness for noise-free images, underpinning its application to enhance feature representation in contexts devoid of noise.
4.4 Re-parameterization for Efficient Inference
To achieve an efficient HDR network that meets the stipulated design prerequisites of low computational complexity and streamlined hardware device deployment, we simplify the TCB reparameterization into a single $3\times 3$ convolution after training. Following previous works [6, 5, 44], we leverage the additivity and homogeneity of convolutions, and we merge the $1\times 1$ extending and $3\times 3$ squeezing convolution into a single $3\times 3$ convolution. Additionally, we combine the Sobel and Laplacian operators into a special $3\times 3$ convolution with a fixed convolution kernel. The $1\times 1$ convolution is achieved by padding the convolution kernel with zeros. As a result, TCB can be transformed into a $3\times 3$ convolution for efficient implementation during the inference stage, as shown in Fig. 3(b). By utilizing TCB, we achieve superior HDR results with improved efficiency.
Figure 5: An illustrative sample of data construction for the proposed alignment-free and motion-aware short-exposure-first selection loss.
@@ 4.4 Re-parameterization for Efficient Inference
Table 2: Performance comparison of different HDR models on our synthetic dataset. #Param and FLOPs represent the total number of network parameters and floating-point operations. The FLOPs and Run Times results are measured on an RTX 3090 device with a resolution of $4096\times 2952$ raw images. Metrics with $\uparrow$ and $\downarrow$ denote higher better and lower better, respectively. The best and second-best performances are in bold and underlined, respectively. "-" indicates the result is not available.
Methods FLOPs #Param Run Time All-Exposure Ratio=4 Ratio=8 Ratio=16
@@
4.5 Loss Functions
Alignment-free and motion-aware short-exposure-first selection loss. In fused-based HDR methods, eliminating ghosting caused by motion inconsistencies between short- and long-exposure pairs is one of the most challenging issues. Previous work [14, 34] commonly employs optical flow, attention mechanisms, and other methods to establish pixel correspondences between short- and long-exposure images. The objective is to suppress ghosting by designing more elaborate fusion strategies. However, these motion estimation and alignment methods are often the most computationally intensive components and cannot be accommodated by the current level of hardware design. On the other hand, unlike other image alignment tasks, such as video motion estimation and stereo matching, short- and long-exposure image fusion in HDR reconstruction does not necessarily require pixel-level correspondence. The reason is that it is challenging to recover sharp object edges due to motion blur. In contrast, short-exposure images exhibit less motion blur distortion. Therefore, ghost artifacts can be suppressed by simply detecting motion regions in long-frame images through some mechanism, discarding these pixels during the fusion process, and relying solely on the corresponding regions in short-exposure images as the exclusive information source for fusion. Based on the same consideration, overexposed regions in the long-exposure image should likewise be discarded in the fusion process. Short-exposure images are often used as reference images in engineering applications.
Based on the above analysis, we devise the strategies for dual-exposure HDR fusion: Firstly, in scenarios involving motion or overexposure within the fused region, we prefer to select short-exposure image that contains more information; Secondly, when the SNR of the short frame is too low, we prefer to select the long-exposure image that contains more information; Thirdly, our strategy is inclined to address ghost artifacts with a higher priority than lower SNR when ghost artifacts and lower SNR concurrently exist.
For the designed fusion strategy, we introduce a plug-and-play alignment-free and motion-aware short-exposure-first selection loss to mitigate the ghost artifacts. We first construct a mask $M$ in the training pairs {$I^{{}^{\prime}}_{l}$, $I^{{}^{\prime}}_{s}$}. Specifically, as shown in Fig. 5, for each patch of an image, we randomly select a rectangle of random length and width from the long-exposure image and then move and overlap it to a random location in the range of -30 to 30 relative to the patch. The patch regions before and after the movement are labeled as 1s in the mask $M$. By introducing the mask $M$ to guide the network to focus on moving and overexposed regions, the model effectively prioritizes the short-exposure information over the long-exposed counterpart within these regions. The masks are used only in training, and the inference stage inputs only short and long exposure frames. The alignment-free and motion-aware short-exposure-first selection loss is denoted as:
$L_{\textrm{AMSS}}=1-\text{MS-SSIM}(\widetilde{I}^{out}\odot M,{I}^{gt}\odot M),$ (15)
where $\odot$ denotes the point-wise multiplication, and $M$ is a binary mask with “1” for motion regions in long frames, and “0” otherwise, MS-SSIM denotes multi-scale structural similarity function. Given a mask $M$ indicating motion and overexposed regions, the above loss formula implements a strategy that encourages short-frame prioritization.
Reconstruction loss. For saturated areas, the L2 loss punishes any deviation of pixel values from the ground truth. This allows the model to select short-exposure information in over-exposed areas.
$L_{\textrm{pix}}=\|\widetilde{I}^{out}-{I}^{gt}\|_{2},$ (16)
To achieve the best HDR reconstruction results, we employ the multi-scale structural similarity loss function loss guide model, which learns short-exposure image information for global motion. By combining these loss functions, our model effectively produces superior results for both areas with motion and saturated regions.
$L_{\textrm{ssim}}=1-\text{MS-SSIM}(\widetilde{I}^{out},{I}^{gt})$ (17)